# Deep Class-Incremental Learning: A Survey

### v1

@@ Abstract
Deep models, e.g., CNNs and Vision Transformers, have achieved impressive achievements in many vision tasks in the closed world. However, novel classes emerge from time to time in our ever-changing world, requiring a learning system to acquire new knowledge continually. For example, a robot needs to understand new instructions, and an opinion monitoring system should analyze emerging topics every day. Class-Incremental Learning (CIL) enables the learner to incorporate the knowledge of new classes incrementally and build a universal classifier among all seen classes. Correspondingly, when directly training the model with new class instances, a fatal problem occurs — the model tends to catastrophically forget the characteristics of former ones, and its performance drastically degrades. There have been numerous efforts to tackle catastrophic forgetting in the machine learning community. In this paper, we survey comprehensively recent advances in deep class-incremental learning and summarize these methods from three aspects, i.e., data-centric, model-centric, and algorithm-centric. We also provide a rigorous and unified evaluation of 16 methods in benchmark image classification tasks to find out the characteristics of different algorithms empirically. Furthermore, we notice that the current comparison protocol ignores the influence of memory budget in model storage, which may result in unfair comparison and biased results. Hence, we advocate fair comparison by aligning the memory budget in evaluation, as well as several memory-agnostic performance measures. The source code to reproduce these evaluations is available at https://github.com/zhoudw-zdw/CIL_Survey/.
@@ 1 Introduction
Recent years have witnessed the rapid progress of deep learning, where deep neural networks have achieved or even surpassed human-level performances in many fields [1, 2, 3]. The typical training process of a deep network requires pre-collected datasets in advance, e.g., large-scale images [4] or texts [5] — the network undergoes training process of the pre-collected dataset multiple epochs. However, training data is often with stream format in the open world [6, 7]. These streaming data cannot be held for long due to storage constraints [8, 9] or privacy issues [10, 11], requiring the model to be updated incrementally with only new class instances. Such requirements trigger the prosperity of the Class-Incremental Learning (CIL) field, aiming to continually build a holistic classifier among all seen classes. The fatal problem in CIL is called catastrophic forgetting, i.e., directly optimizing the network with new classes will erase the knowledge of former ones and result in irreversible performance degradation. Hence, how to effectively resist catastrophic forgetting becomes the core problem in building CIL models.
@@
Figure 1: The setting of class-incremental learning. Non-overlapping classes arrive sequentially, and the model needs to learn to classify all the classes incrementally. After the learning process of each task, the model is evaluated among all seen classes. An ideal model should perform well in the newly learned classes and remember the former without catastrophic forgetting.
@@
Figure 2: The setting of Class-Incremental Learning (CIL), Task-Incremental Learning (TIL), and Domain-Incremental Learning (DIL). CIL and TIL share the same training protocol, while TIL is much easier during inference, i.e., only requiring classifying among corresponding label spaces. DIL refers to the data stream with distribution change, where new tasks contain the same classes from different domains, e.g., cartoon and clip-art.
@@
Apart from class-incremental learning, there are other fine-grained settings addressing the incremental learning problem, e.g., Task-Incremental Learning (TIL) and Domain-Incremental Learning (DIL). We show these three protocols in Figure 2. TIL is a similar setting to CIL, and both of them observe incoming new classes in new tasks. However, the difference lies in the inference stage, where CIL requires the model to differentiate among all classes. By contrast, TIL only requires classifying the instance among the corresponding task space. In other words, it does not require cross-task discrimination ability. Hence, TIL is easier than CIL, which can be seen as a particular case of CIL. On the other hand, DIL concentrates on the scenario with concept drift or distribution change [13, 14], where new tasks contain instances from different domains but with the same label space. In this case, new domains correspond to the images in clip-art format. In this paper, we concentrate on the CIL setting, which is a more challenging scenario in the open world.
@@
There is also research about CIL before the prosperity of deep learning [15]. Typical methods try to solve the catastrophic forgetting problem with traditional machine learning models. However, most of them address the incremental learning within two tasks, i.e., the model is only updated with a single new stage [16, 17, 18]. Furthermore, the rapid development of data collection and processing requires the model to grasp long-term and massive-scale data streams that traditional machine learning models cannot handle. Correspondingly, deep neural networks with powerful representation ability well suit these requirements. As a result, deep learning-based CIL is becoming a hot topic in the machine learning and computer vision community.
@@
There are several related surveys discussing the incremental learning problem. For example, [11] focuses on the task-incremental learning problem and provides a comprehensive survey. [19] is a related survey on the class-incremental learning field, while it only discusses and evaluates the methods till 2020. However, with the rapid development of the CIL field, many great works are emerging day by day, which substantially boost the performance of benchmark settings [20, 21, 22, 23]. On the other hand, with the prosperity of Vision Transformer (ViT) [24] and pre-trained models, a heated discussion about ViT in CIL is attracting the attention of the community. Other surveys either focuses on the specific field [25, 26, 27] or lacking the performance evolution among state-of-the-arts [28, 29, 30]. Hence, it is urgent to provide an up-to-date survey containing popular methods to speed up the development of the CIL field.
@@
In this paper, we aim for a comprehensive review of deep class-incremental learning methods from three aspects, i.e., data-centric, model-centric, and algorithm-centric. We also provide a holistic comparison among different kinds of methods over benchmark datasets, i.e., CIFAR100 [31] and ImageNet100/1000 [4]. On the other hand, we highlight an important factor in CIL model evaluation, i.e., memory budget, and advocate fair comparison among different methods with an aligned budget. Correspondingly, we holistically evaluate the extensibility of CIL models with budget-agnostic measures. In general, the contribution of this survey can be summarized as follows:
  • We provide a comprehensive survey of deep CIL, including problem definitions, benchmark datasets, and different families of CIL methods. We organize these algorithms taxonomically (Table I) and chronologically (Figure 3) to give a holistic overview of state-of-the-art.
  • We provide a rigorous and unified comparison among different methods on several publicly available datasets, including traditional CNN-backed and modern ViT-backed methods. We also discuss the insights and summarize the common rules to inspire future research.
@@
  • To boost real-world applications, CIL models should be deployed not only on high-performance computers but also on edge devices. Therefore, we advocate evaluating different methods holistically by emphasizing the effect of memory budgets. Correspondingly, we provide a comprehensive evaluation of different methods given specific budgets as well as several new performance measures.
The rest of this paper is organized as follows. First, we give the problem definition of class-incremental learning in Section 2. Afterward, we organize current CIL methods taxonomically in Section 3 and provide comprehensive evaluations in Section 4. Finally, we summarize future directions of class-incremental learning in Section 5 and then conclude this paper.
@@ 2 Preliminaries
In this section, we first give the problem formulation of class-incremental learning. We also discuss the other variations and CIL model decomposition. Afterward, we introduce a commonly used auxiliary set, i.e., exemplar set for CIL.
@@ 2.1 Problem Formulation
Class-Incremental Learning aims to learn from an evolutive stream with incoming new classes [32]. Assume there is a sequence of $B$ training tasks [^1] Also known as ‘session,’ ‘phase,’ and ‘stage.’ $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$ without overlapping classes, where $\mathcal{D}^{b}=\left\{\left({\bf x}_{i}^{b},y_{i}^{b}\right)\right\}_{i=1}^{n_{b}}$ is the $b$-th incremental step with $n_{b}$ training instances. ${\bf x}_{i}^{b}\in\mathbb{R}^{D}$ is an instance of class $y_{i}\in Y_{b}$, $Y_{b}$ is the label space of task $b$, where $Y_{b}\cap Y_{b^{\prime}}=\varnothing$ for $b\neq b^{\prime}$. We can only access data from $\mathcal{D}^{b}$ when training task $b$. The ultimate goal of CIL is to continually build a classification model for all classes. In other words, the model should not only acquire the knowledge from the current task $\mathcal{D}^{b}$ but also preserve the knowledge from former tasks. After each task, the trained model is evaluated over all seen classes $\mathcal{Y}_{b}=Y_{1}\cup\cdots Y_{b}$. Formally, CIL aims to fit a model $f({\bf x}):X\rightarrow\mathcal{Y}_{b}$, which minimizes the expected risk:
$f^{*}=\underset{f\in\mathcal{H}}{\operatorname{argmin}}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{t}^{1}\cup\cdots\mathcal{D}_{t}^{b}}\mathbb{I}(y\neq f(\mathbf{x}))\,$, (1)
@@
Class Overlapping: Typical CIL setting assumes $Y_{b}\cap Y_{b^{\prime}}=\varnothing$ for $b\neq b^{\prime}$, i.e., there are no overlapping classes in different tasks. However, in the real world, it is common to observe the old classes emerging in new tasks [33]. When $Y_{b}\cap Y_{b^{\prime}}\neq\varnothing$, the setting is called blurry class-incremental learning (Blurry CIL) [34, 35, 36]. It enables the model to revisit former instances in the later stage, which weakens the learning difficulty. Hence, this paper focuses on the typical CIL setting with no overlapping classes.
@@
Online CIL: Although data comes with stream format in CIL, the model can conduct multi-epoch training with each training task $\mathcal{D}^{b}$, i.e., offline training within each task. There are some works addressing fully online (one-pass) class-incremental learning, where each batch can be processed once and then dropped [37, 38]. It is a specific case of the current setting, and we concentrate on the generalized CIL setting in this paper.
Table I: The taxonomy of class-incremental learning. We classify CIL methods into data-centric, model-centric, and algorithm-centric methods and further decompose them into six sub-categories. The shade color in the last column denotes the subcategory, which is consistent with Figure 3.
Algorithm Category Subcategory Reference
$\S$ 3.1 Data-Centric Class-Incremental Learning $\S\penalty 10000\$3.1.1 Data Replay Direct Replay [39, 35, 40, 41, 42, 43, 44, 45, 46, 47]
Generative Replay [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]
$\S\penalty 10000\$3.1.2: Data Regularization [59, 60, 40, 61, 62, 63]
$\S$ 3.2 Model-Centric Class-Incremental Learning $\S\penalty 10000\$3.2.1 Dynamic Networks Neuron Expansion [64, 65, 66]
Backbone Expansion [67, 68, 69, 70, 71, 20, 21, 22]
Prompt Expansion [23, 72, 73, 74, 75, 76]
$\S\penalty 10000\$3.2.2: Parameter Regularization [77, 78, 39, 79, 80, 81, 82, 83, 84]
$\S$ 3.3 Algorithm-Centric Class-Incremental Learning $\S\penalty 10000\$3.3.1 Knowledge Distillation Logit Distillation [85, 86, 32, 87, 88, 89, 90, 91, 92]
Feature Distillation [93, 94, 95, 96, 97, 98, 99, 100]
Relational Distillation [101, 102, 103, 104, 105]
$\S\penalty 10000\$3.3.2 Model Rectify Feature Rectify [106, 107, 108, 109, 110, 111]
Logit Rectify [93, 112, 87, 113, 114]
Weight Rectify [115, 116, 117]
@@ 2.2 Exemplars and Exemplar Set
As defined in Definition 1, in each incremental task, the model is only able to access the current dataset $\mathcal{D}^{b}$. This helps to preserve user privacy and release the storage burden. However, this restriction is relaxed in some cases, and the model can keep a relatively small set, namely exemplar set, to reserve the representative instances from former tasks.
@@
Exemplar Set [^3] Also known as the ‘replay buffer’ and ‘memory buffer.’ We use them interchangeably in this paper. is an extra collection of instances from former tasks $\mathcal{E}=\left\{\left({\bf x}_{j},y_{j}\right)\right\}_{j=1}^{M}$, $y_{j}\in\mathcal{Y}_{b-1}$. With the help of the exemplar set, the model can utilize $\mathcal{E}\cup\mathcal{D}^{b}$ for the update within each task. The model manages the exemplar set after the training process of each task.
@@
Exemplar Selection: Exemplars are representative instances of each known class, which need to be selected from the entire training set. An intuitive way to choose the exemplars is to randomly sample exemplars for each class, which results in diverse instances. By contrast, a commonly used strategy is called herding [118, 32], aiming to select the most representative ones of each class. Given the instance set $X=\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}$ from class $y$, herding first calculates the class center with current embedding $\phi(\cdot)$:
$\mu_{y}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\phi({\bf x}_{i})\,.$ (3)
Afterward, it calculates and ranks the distance of each instance to the class center $\left\|\mu_{y}-\phi({\bf x}_{i})\right\|$ in ascending order. The exemplars are then selected based on the ranking, e.g., the top-$\frac{M}{|\mathcal{Y}_{b}|}]$ instances with the least distance. Since the class center can be seen as the most representative pattern of each class, selecting exemplars near the class center also enhances the representativeness of exemplars. Herding is now a commonly-used strategy to select exemplars in CIL, and we also adopt it in this paper.

### v2

# Class-Incremental Learning: A Survey

@@ Abstract
Deep models, e.g., CNNs and Vision Transformers, have achieved impressive achievements in many vision tasks in the closed world. However, novel classes emerge from time to time in our ever-changing world, requiring a learning system to acquire new knowledge continually. Class-Incremental Learning (CIL) enables the learner to incorporate the knowledge of new classes incrementally and build a universal classifier among all seen classes. Correspondingly, when directly training the model with new class instances, a fatal problem occurs — the model tends to catastrophically forget the characteristics of former ones, and its performance drastically degrades. There have been numerous efforts to tackle catastrophic forgetting in the machine learning community. In this paper, we survey comprehensively recent advances in class-incremental learning and summarize these methods from several aspects. We also provide a rigorous and unified evaluation of 17 methods in benchmark image classification tasks to find out the characteristics of different algorithms empirically. Furthermore, we notice that the current comparison protocol ignores the influence of memory budget in model storage, which may result in unfair comparison and biased results. Hence, we advocate fair comparison by aligning the memory budget in evaluation, as well as several memory-agnostic performance measures. The source code is available at https://github.com/zhoudw-zdw/CIL_Survey/.
@@ 1 Introduction
Recent years have witnessed the rapid progress of deep learning, where neural networks have achieved or even surpassed human-level performances in many fields [1, 2, 3]. The typical training process of a deep network requires pre-collected datasets in advance, e.g., large-scale images or texts — the network undergoes the training process of the pre-collected dataset multiple epochs. However, training data is often with stream format in the open world [4]. These streaming data cannot be held for long due to storage constraints [5] or privacy issues [6], requiring the model to be updated incrementally with only new class instances. Such requirements trigger the prosperity of the Class-Incremental Learning (CIL) field, aiming to continually build a holistic classifier among all seen classes. The fatal problem in CIL is called catastrophic forgetting, i.e., directly optimizing the network with new classes will erase the knowledge of former ones and result in irreversible performance degradation. Hence, how to effectively resist catastrophic forgetting becomes the core problem in building CIL models.
@@
Figure 1: The setting of CIL. Non-overlapping classes arrive sequentially, and the model needs to learn to classify all the classes incrementally. After learning each task, the model is evaluated among all seen classes. An ideal model should perform well in the newly learned classes and remember the former without forgetting.
@@
Figure 2: The setting of Class-Incremental Learning (CIL), Task-Incremental Learning (TIL), and Domain-Incremental Learning (DIL). CIL and TIL share the same training protocol, while TIL is much easier during inference, i.e., only requiring classifying among corresponding label spaces. DIL refers to the data stream with distribution change, where new tasks contain the same classes from different domains, e.g., cartoon and clip-art. The distinction of these scenarios is proposed by [8].
@@
Since the data stream comes continually and requires training a lifelong time, incremental learning is also known as ‘continual learning[6] or ‘lifelong learning[9]. We interchangeably use these concepts in this paper. Apart from class-incremental learning, there are other fine-grained settings addressing the incremental learning problem, e.g., Task-Incremental Learning (TIL) and Domain-Incremental Learning (DIL) [8]. We show these three protocols in Figure 2. TIL is a similar setting to CIL, and both of them observe incoming new classes in new tasks. However, the difference lies in the inference stage, where CIL requires the model to differentiate among all classes. By contrast, TIL only requires classifying the instance among the corresponding task space. In other words, it does not require cross-task discrimination ability. Hence, TIL is easier than CIL, which can be seen as a particular case of CIL. On the other hand, DIL concentrates on the scenario with concept drift or distribution change [10], where new tasks contain instances from different domains but with the same label space. In this case, new domains correspond to the images in clip-art format. In this paper, we concentrate on the CIL setting, which is a more challenging scenario in the open world.
@@
There is also research about CIL before the prosperity of deep learning [11]. Typical methods try to solve the catastrophic forgetting problem with traditional machine learning models. However, most of them address the incremental learning within two tasks, i.e., the model is only updated with a single new stage [12, 13]. Furthermore, the rapid development of data collection and processing requires the model to grasp long-term and massive-scale data streams that traditional machine learning models cannot handle. Correspondingly, deep neural networks with powerful representation ability well suit these requirements. As a result, deep learning-based CIL is becoming a hot topic in the machine learning and computer vision community.
@@
There are several related surveys discussing the incremental learning problem. For example, [6] focuses on the task-incremental learning problem and provides a comprehensive survey. [14] is a related survey on the class-incremental learning field, while it only discusses and evaluates the methods till 2020. However, with the rapid development of the CIL field, many great works are emerging day by day, which substantially boost the performance of benchmark settings [15, 16, 17, 18]. On the other hand, with the prosperity of Vision Transformer (ViT) [19] and pre-trained models, a heated discussion about ViT in CIL is attracting the attention of the community. Other surveys either focus on the specific field [20, 21] or lack the performance evolution among state-of-the-arts [8, 22, 23]. Hence, it is urgent to provide an up-to-date survey containing popular methods to speed up the development of the CIL field.
@@
In this paper, we aim for a comprehensive review of class-incremental learning methods and divide them into seven categories. We also provide a holistic comparison among different kinds of methods over benchmark datasets, i.e., CIFAR100 [24] and ImageNet100/1000 [25]. On the other hand, we highlight an important factor in CIL model evaluation, i.e., memory budget, and advocate fair comparison among different methods with an aligned budget. Correspondingly, we holistically evaluate the extensibility of CIL models with budget-agnostic measures.
@@
In general, the contribution of this survey can be summarized as follows: 1): We provide a comprehensive survey of CIL, including problem definitions, benchmark datasets, and different families of CIL methods. We organize these algorithms taxonomically (Table I) and chronologically (Figure 3) to give a holistic overview of state-of-the-art. 2): We provide a rigorous and unified comparison among different methods on several publicly available datasets, including traditional CNN-backed and modern ViT-backed methods. We also discuss the insights and summarize the common rules to inspire future research. 3): To boost real-world applications, CIL models should be deployed not only on high-performance computers but also on edge devices. Therefore, we advocate evaluating different methods holistically by emphasizing the effect of memory budgets. Correspondingly, we provide a comprehensive evaluation of different methods given specific budgets as well as several new performance measures.
@@ 2.1 Problem Formulation
Class-Incremental Learning aims to learn from an evolutive stream with new classes. Assume there is a sequence of $B$ training tasks [^1] Also known as ‘session,’ ‘phase,’ and ‘stage.’ $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$ without overlapping classes, where $\mathcal{D}^{b}=\left\{\left({\bf x}_{i}^{b},y_{i}^{b}\right)\right\}_{i=1}^{n_{b}}$ is the $b$-th incremental step with $n_{b}$ training instances. ${\bf x}_{i}^{b}\in\mathbb{R}^{D}$ is an instance of class $y_{i}^{b}\in Y_{b}$, $Y_{b}$ is the label space of task $b$, where $Y_{b}\cap Y_{b^{\prime}}=\varnothing$ for $b\neq b^{\prime}$. We can only access data from $\mathcal{D}^{b}$ when training task $b$. The ultimate goal of CIL is to continually build a classification model for all classes. In other words, the model should not only acquire the knowledge from the current task $\mathcal{D}^{b}$ but also preserve the knowledge from former tasks. After each task, the trained model is evaluated over all seen classes $\mathcal{Y}_{b}=Y_{1}\cup\cdots Y_{b}$. Formally, CIL aims to fit a model $f({\bf x}):X\rightarrow\mathcal{Y}_{b}$, which minimizes the expected risk:
$f^{*}=\underset{f\in\mathcal{H}}{\operatorname{argmin}}\kern 5.0pt\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{t}^{1}\cup\cdots\mathcal{D}_{t}^{b}}\mathbb{I}(y\neq f(\mathbf{x}))\,$, (1)
@@
Class Overlapping: Typical CIL setting assumes $Y_{b}\cap Y_{b^{\prime}}=\varnothing$ for $b\neq b^{\prime}$, i.e., there are no overlapping classes in different tasks. However, in the real world, it is common to observe the old classes emerging in new tasks. When $Y_{b}\cap Y_{b^{\prime}}\neq\varnothing$, the setting is called blurry class-incremental learning (Blurry CIL) [26]. It enables the model to revisit former instances in the later stage, which weakens the learning difficulty.
Number of Instances: Typical CIL setting assumes the training data of each class is balanced and many-shot (e.g., hundreds or thousands). However, the data collection may face challenges in the real world, e.g., we can only collect a limited number of training instances for rare birds. When the training instances of new classes are limited (e.g., 5-shot per class), the setting is called Few-Shot Class-Incremental Learning (FSCIL) [27]. When the training instances are highly imbalanced and long-tailed, the setting is called Long-Tailed Class-Incremental Learning (LTCIL) [28].
Online CIL: Although data comes with stream format, the model can conduct multi-epoch training with each task, i.e., offline training within each task. There are some works addressing fully online (one-pass) CIL, where each batch can be processed once and then dropped [29]. It is a specific case of the current setting, and we concentrate on the generalized CIL setting in this paper.
@@
Backbones: As defined in Eq. 2, the predictions are derived by feeding instance $\bf x$} into the embedding function and linear layer. The embedding function is designed to project input instances into the embedding space to reflect its semantic information. Hence, ideal CIL algorithms should work for any type of backbones, e.g., multi-layer perceptron (MLP) [30], convolutional neural network (CNN) [31], and Vision Transformer (ViT) [19]. Specifically, we often treat the whole image as the input of MLP and CNN and utilize the final product as the embedding. By contrast, ViT transforms the image into a sequence of patches for patch features. These patches are then fed forward to the self-attention modules [32] and MLP layer to get the contextualized information. Typical ViT appends an extra token (i.e., [CLS] token) to the set of patches and utilizes the final representation of [CLS] token as the embedding. Since ViT relies on the self-attention mechanism to relate and adjust patch-wise features, it is intuitive to influence the embedding context by adding task-specific tokens as the input. The basic difference between these backbones triggers different tuning algorithms in CIL, which will be further discussed in Section 3.3.
Baseline in CIL: In CIL, the typical baseline is sequential finetuning the model with the current dataset $\mathcal{D}^{b}$ (denoted as ‘Finetune’), whose loss function can be denoted as:
$\mathcal{L}=\sum_{({\bf x},y)\in\mathcal{D}^{b}}\ell(f({\bf x}),y)\,$, (3)
where $\ell(\cdot,\cdot)$ measures the discrepancy between inputs, e.g., cross-entropy loss. Finetune is known as the baseline method for CIL since it only concentrates on learning the new concepts in the current task. Consequently, it suffers severe forgetting since the model pays no attention to former ones.
Table I: The taxonomy of CIL. The shade color in the last column denotes the subcategory, which is consistent with Figure 3.
Algorithm Category Reference
$\S\penalty 10000\$3.1 Data Replay
Direct Replay
[33, 26, 34, 35, 36, 37, 38, 39, 40]
Generative Replay
[41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]
$\S\penalty 10000\$3.2: Data Regularization
[53, 54, 34, 55, 56, 57, 58]
$\S\penalty 10000\$3.3 Dynamic Networks
Neuron Expansion
[59, 60, 61, 45]
Backbone Expansion
[62, 63, 64, 65, 66, 67, 15, 16, 17, 68]
Prompt Expansion
[18, 69, 70, 71, 72]
$\S\penalty 10000\$3.4: Parameter Regularization
[73, 74, 33, 75, 76, 77, 78, 79, 80]
$\S\penalty 10000\$3.5 Knowledge Distillation
Logit Distillation
[81, 82, 83, 84, 85, 86, 87, 88]
Feature Distillation
[89, 90, 91, 92, 93, 94, 95, 96, 97, 98]
Relational Distillation
[99, 100, 101, 27, 102, 103]
$\S\penalty 10000\$3.6 Model Rectify
Feature Rectify
[104, 105, 106, 107, 108, 109]
Logit Rectify
[89, 110, 83, 111]
Weight Rectify
[112, 113, 114]
$\S\penalty 10000\$3.7: Template-Based Classification
[82, 37, 115, 104, 116, 117, 118, 119, 120, 121, 122, 123]
@@ 2.2 Exemplars and Exemplar Set
As defined in Definition 1, in each incremental task, the model can only access the current dataset $\mathcal{D}^{b}$. This helps to preserve user privacy and release the storage burden. However, this restriction is relaxed in some cases, and the model can keep a relatively small set, namely exemplar set, to reserve the representative instances from former tasks.
@@
Exemplar Set [^3] Also known as the ‘replay buffer’ and ‘memory buffer.’ is an extra collection of instances from former tasks $\mathcal{E}=\left\{\left({\bf x}_{j},y_{j}\right)\right\}_{j=1}^{M}$, $y_{j}\in\mathcal{Y}_{b-1}$. With the help of the exemplar set, the model can utilize $\mathcal{E}\cup\mathcal{D}^{b}$ for the update within each task. The model manages the exemplar set after the training process of each task.
@@
Exemplar Selection: Exemplars are representative instances of each class, which need to be selected from the entire training set. An intuitive way to choose the exemplars is random sampling, which results in diverse instances. By contrast, a commonly used strategy is called herding [124, 82], aiming to select the most representative ones of each class. Given the instance set $X=\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}$ from class $y$, herding first calculates the class center with current embedding $\phi(\cdot)$: ${\bm{\mu}}_{y}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\phi({\bf x}_{i})$. Afterward, it iteratively appends instances into the exemplar set:
${\bm{p}}_{k}\leftarrow\underset{{\bf x}\in X}{\operatorname{argmin}}\left\|{\bm{\mu}}_{y}-\frac{1}{k}\left[\phi({\bf x})+\sum_{j=1}^{k-1}\phi\left({\bm{p}}_{j}\right)\right]\right\|\,$, (4)
until $k$ reaches the memory bound of each class ($\frac{M}{|\mathcal{Y}_{b}|}]$). The exemplar set of class $y$ is the concatenation of $\{{\bm{p}}_{1},{\bm{p}}_{2},\cdots{\bm{p}}_{[\frac{M}{|\mathcal{Y}_{b}|}]}\}$. Eq. 4 ensures that the average feature vector over all exemplars selected thus far is closest to the class mean. Since the class center can be seen as the most representative pattern of each class, selecting exemplars near the center also enhances the representativeness. Herding is now a commonly used strategy to select exemplars in CIL, and we also adopt it in this paper.